Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
berni - 22:56 Thursday 03 April 2025 (66500) Print this report
Operator Report - Afternoon shift

ITF found in Commissioning mode with locked arms; calibration activity in progress.

The calibration activity was concluded around 14:10 UTC; after that a strong earthquake (M 6.9 - Reykjanes Ridge)  prevented to relock.

At 16:16 UTC the ITF reached LN3; then we contacted Valerio to perform the WE kick. Unfurtunately the ITF unlocked after 1s.

A second kick was scheduled at 20:00 UTC but Valerio was not available; I did not understand if he will perform the kick later in the night...

For this reason I left the ITF in Commissioning mode with Autorelock_failsafe engaged.

 

SBE

SPRB opened because of the earthquake; properly closed.

Images attached to this report
Comments to this report:
rolland - 21:54 Thursday 03 April 2025 (66507) Print this report

WE PCal switched off today around 17h30 UTC. The laser driver has been switched off to prevent any restart of the PCal during the WE mirror replacement.

The DMS flags associated to the WE Pcal have been shelved  for two weeks  (the PCal ones around 17h30, the hrec bias ones based on the WE PCal around 19h50).

Search Help
×

Warning

×